OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (luigi0210):

Take the slope of the equation, which is 3, then use the coordinates given. Plug it all into y1-y=m(x-x1)

OpenStudy (luigi0210):

so y+4=3(x-2)

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

don't forget about that 4

OpenStudy (anonymous):

oh, so it's A?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Yes

OpenStudy (anonymous):

awesome thanks homie